Question 4 (1 point) Saved You are bearish on Telecom and decide to sell short 100 shares at the current market price of $50 per share. How much in cash or securities must you put into your brokerage account if the brokers initial margin requirement is 50% of the value of the short position? Enter your answer without a dollar sign Your Answer: 2500 Answer Question 5 (1 point) (continued from the previous question) How high can the price of the stock go before you get a margin call if the maintenance margin is 40% of the value of the short position? Round your answer to two decimal places and enter it without dollar sign Your Answer: Answer
Answer #1

Current Securities or Cash in the Trading Account = $2500

Number of Shares = 100

At 40% maintenance margin value of short position for 100 shares with trading amount of $2500:

= 2500/0.40 = $6250

Hence, Price of stock at 40% maintenance margin call = 6250/100 = $62.5